A wire of mass 0.3gm is lying horizontal on water surface.surface tension is 70ร—10^-3N/m.ย  The maximum length of wire so that it may not sink is... . .... . .. .ANS: 2.1cmย 

Grade:12th pass

1 Answers

Darshini Prakash
5 years ago
2S = F/L ; F=mg
(Where S=Surface tension; F=Force acting on the body; L=Length of the body)
Substituting the given values we get:
2(70*10-3ย ) = (0.3*10-3)*10/Lย  ย  ย  ย (**WE SHOULD SUBSTITUTE ALL THE VALUES IN SAME UNITS**)
L=0.021mย  orย  L=2.1cm
ย